### v1

# Rate of epidemic spreading on complex networks

Abstract
The initial phase of an epidemic is often characterised by an exponential increase in the number of infected individuals. In well-mixed populations, this exponential increase is controlled by the basic reproduction number $R_{0}$ and the distribution of times between consecutive infection along an infection chain. However, we are still lacking a general understanding of how epidemics spread when individual interactions form a complex network. Here, we derive an expression for the rate of exponential spread of an epidemic on a complex network. We find that this rate is affected by the degree distribution, the network assortativity, and the level of clustering. Our result holds for a broad range of networks, aside from networks with very broad degree distribution, where no clear exponential regime is present. The theory presented in this paper bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
Epidemic modelling permits to rationalise evidences, test hypotheses, and inform disease control and other policy making. Epidemiological models should incorporate the key features that determine the epidemic spreading. One such feature is the fact that individuals are arranged in social structures [1, 2]. These social structures play a crucial role in the spreading of an epidemic, and in the effectiveness of intervention strategies [3, 4]. In these social structures, the number of contacts per individual is highly variable. For example, at the early stages of COVID-19, the individuals with most contacts contributed disproportionately to the epidemic [5, 6].
A convenient way of modelling social structures is by means of complex networks [7], in which nodes represent individuals and links signify that pairs of individuals are in close contact. A success of network theory has been to determine the epidemic threshold on networks [8], i.e., the transition point from a localised outbreak to a widespread epidemic. This theory has shown that networks whose degree distribution has a diverging variance due to heavy tails [9, 10] are extremely vulnerable to epidemics. A paradigmatic example is found in the Barabási-Albert model [11], where a network is constructed by preferentially attaching new nodes to existing nodes with high degree. Another property of complex networks which is relevant for epidemic spreading is clustering, which is a measure of how commonly nodes tend to form tightly connected communities. Many networks in nature possess both fat tails in their degree distribution and high level of clustering [12].
A second key feature is the fact that the infectiousness of an individual, i.e., the rate at which an infected individuals infects its contacts, tends to change in time in a way that is particular to each disease. The underlying reason is that the infectiousness incorporates multiple evolving factors, such as the viral load of an individual, its social behaviour, and possibly environmental factors [13]. This time-dependence profoundly impacts the way a disease spreads.
Social structures and individual infectiousness are the key factors in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during emergence of novel variants, often requiring rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioural change as a consequence of the epidemic [14], and the possibility of being reinfected.
Despite the vast body of theory on epidemics in networks [10], results on the initial exponential regime are surprisingly sparse. A recent theory estimates the average time it takes for an epidemic to reach an individual at a certain distance from patient zero [15]. The spread of an epidemic on a network can be exactly solved [16], but this calculation becomes practically unfeasible unless the network is very small. Other approaches avoid a detailed description of social structures, and effectively model the heterogeneity of the population by assigning different risk propensities to individuals [17, 18].
In this paper, we directly estimate the exponential rate of epidemic spreading on a complex network. Our approach is based on classic ideas of theoretical epidemiology, that we extend to complex networks.
Epidemic spreading in well-mixed populations
We begin by summarising classic results of epidemic modelling in well-mixed populations [13]. In this case, the cumulative number $I(t)$ of infected individuals at time $t$ is governed by the renewal equation
[…]
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual, and the dot denotes a time derivative. The function $w(\tau)$ is the infectiousness, defined as the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. We assume to be in the exponential regime, $I(t)\sim e^{\Lambda t}$, where $\Lambda$ is the rate of exponential spread. Substituting this assumption into Eq. (1) we obtain the Euler-Lotka equation
$\frac{1}{R_{0}}=\langle e^{-\Lambda\tau}\rangle\,,$ (2)
[…]
Epidemic spreading in networks.
[…]
Infected nodes cannot get reinfected and are never formally considered as recovered, although their infectiousness may wane over time. Our model can therefore be seen as a Susceptible-Infected model on a network with non-Markovian infection times. The Markov property only holds in the limiting case of constant $\lambda(\tau)$, and consequently exponentially distributed $\psi(\tau)$. This limiting case is often unrealistic. For example, the infectiousness of COVID-19 markedly increases for 3–6 days and then decay for about 11 days [19], and is therefore poorly described by an exponential distribution.
In the course of the epidemic, the number of infected rapidly grows before saturating (Fig. 1c). An alternative to the temporal perspective is to represent the epidemic in generations (Fig. 1d). We call $Z_{n}$, $n=0,1,\dots$ the average number of infected individuals in the $n$th generation. The initial generation, $n=0$, is constituted by the patient zero, $Z_{0}=1$. In this representation, the infected individuals form a tree (Fig. 1e).
Figure 1: Epidemic spreading on complex networks. a: Epidemic on a network. Nodes can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbours, to whom they are connected via links. Individuals may have different numbers of neighbours. Due to the presence of cycles, some neighbours may have already been infected by someone else, so that the transmission does not result in an infection. b: Infected individuals spread the disease at a time-dependent rate $\lambda(\tau)$, where $\tau$ is the time since they were infected. The resulting infection times are distributed according to $\psi(\tau)$. c: Example of an epidemic trajectory starting with one infected individual. After an outbreak, the epidemic spreads at an exponential spreading rate $\Lambda$. The exponential phase ends in a saturation phase, where a sizeable fraction of the individuals are already infected. d: Epidemic tree. The epidemic starts with a first infected individual (patient zero) and resulting infections are represented by continuous lines. Dashed lines represents different generations.
Reproduction number on a complex network
We define the basic reproduction number $R_{0}$ for a network as the average number of contacts of an infected individual in the early stage of the epidemic, excluding the one from which the infection came from. We also define the effective reproduction number $R$ as the average number of infections caused by an infected individual. In contrast with $R_{0}$, $R$ does not include contacts that are infected by other individuals. This difference is crucial in clustered networks.
Our initial goal is to estimate $R$ for a complex network. To this aim, we characterise individuals by their degree $k$. We introduce the reproduction matrix $M_{kk^{\prime}}$, whose elements are the average numbers of susceptible individuals of degree $k$ connected to a node of degree $k^{\prime}$. The reproduction matrix is expressed by
$M_{kk^{\prime}}=(k^{\prime}-1-m_{kk^{\prime}})P(k|k^{\prime})$ (3)
where $P(k|k^{\prime})$ is the probability that a neighbour of a node of degree $k^{\prime}$ has degree $k$ and $m_{kk^{\prime}}$ is the average number of triangles a $kk^{\prime}$-link is part of. The term proportional to $m_{kk^{\prime}}$ in Eq. (3) discounts for contacts of a given individual that were already infected by someone else, due to the presence of triangles in the network. In Eq. (3), we neglect higher order loops such as squares. The reproductive matrix given in Eq. (3) has been used to study percolation on networks [20].
On the one hand, assuming that $R$ remains constant between generations, the average number of infected individuals at the $n$th generation is equal to $Z_{n}\sim R^{n}$, up to the leading order. On the other hand, the number of infected individuals $Z_{n}$ scales as the $n$-th power of the reproduction matrix (see Methods). This permits us to identify $R$ as the leading eigenvalue of the reproduction matrix. The associated right eigenvector $v_{k^{\prime}}$ can be interpreted as the degree distribution of the infected individuals.
For large networks, diagonalising the reproduction matrix can be unwieldy. In the simple case where there is no clustering $(m_{kk^{\prime}}=0$ for all links) and the degrees of connected nodes are uncorrelated, we have
$R=R_{0}=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}\,,$ (4)
where the bar represents an average over the degree distribution. Equation (4) is a classic result for epidemic spreading on networks [9, 21, 8]. We seek to extend it to the general case of clustered and correlated networks. We quantify degree correlations by the assortativity coefficient $r$. When $r>0$, nodes with high degree tend to connect together whereas when $r<0$, they tend to connect to lower degree nodes. We find that, for small clustering coefficients and degree correlation, the basic reproduction number $R$ is approximated by
$R_{\mathrm{pert}}=(1-r)(R_{0}-m_{1})+\frac{r}{R_{0}-m_{1}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)\,,$ (5)
where $c_{k}$ is the probability that two neighbours of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$ (see Methods). The reason for calling $R_{\mathrm{pert}}$ the expression in Eq. (5) is that this is obtained by a perturbative expansion in $r$ (see Methods).
We compare the expression for $R_{\mathrm{pert}}$ given in Eq. (5) to the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. We find that the relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the networks we considered, except for Barabási-Albert with $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$). We shall scrutinize the reason for this discrepancy later in the paper. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (5) for a given network, we always interpret the averages as empirical averages. Because of this choice, moments of the degree distribution are always finite, even in very heterogeneous networks. In practical cases, Eq. (5) predicts that the reproduction number $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case [22]. However, the coefficient multiplying $r$ in Eq. (5) is usually rather small, except when the degree distribution has heavy tails. In contrast, clustering systematically lowers $R_{\mathrm{pert}}$, regardless of the degree distribution. Therefore, approximating the reproduction number with $R_{0}$ leads to an overestimate in the presence of clustering.
Table 1: Effect of clustering and assortativity on the reproduction number.
Network Assortativity $r$ Clustering $\overline{c}$ $R_{0}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ $R_{\mathrm{pert}}$ $R$
$\text{BA}_{1}$0.011 0 13.9 13.9 6.6 6.8 8.5
$\text{BA}_{2}$0.008 0.0001 20.4 20.3 14.2 14.4 15.5
$\text{BA}_{12}$0.004 0.0003 85.4 83.6 80.4 79.5 80.8
ER6 0.001 0.000 6.0 6.0 6.0 6.0 6.0
[…]
WS8,2 -0.022 0.334 7.2 4.9 7.0 4.9 4.9
WS6,2 -0.029 0.315 5.2 3.6 5.0 3.6 3.6
$\text{LN}_{4,1}\;\alpha=0.5$ 0.007 0.161 3.7 3.2 3.8 3.2 3.2
Parameters associated with the epidemic spreading for different models of complex networks: BAm: Barabási-Albert model with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbours and rewiring probability $p$. $\text{LN}_{4,1}\;\alpha=0.5$ is generated from a modified version of the configuration model [23] with a Lognormal degree distribution with mean $4$ and variance $1$. This modified algorithm allows us for tuning the clustering coefficients as $c_{k}\propto(k-1)^{-\alpha}$ for a chosen $\alpha>0$. $\text{ER}_{10}\;\alpha=0.2$: A Erdős-Rényi network in which we imposed clustering using the same algorithm. We have defined the average clustering coefficient $\bar{c}=\sum_{k}c_{k}P(k)$. The reproduction number $R$ is the numerically computed leading eigenvalue of Eq. (3), whereas we obtain $R_{\mathrm{pert}}$ by evaluating Eq. (5). To assess the impact of degree correlations and clustering, we also compute $R_{\mathrm{pert}}$ for $\overline{c}=0$.
Network Euler-Lotka equation
We now use our estimate of the reproduction number to predict how epidemics spread. The average total number ${I}_{k}(t)$ of infected individual of degree $k$ at time $t$ is governed by the renewal equation
$\dot{I}_{k}(t)=\sum_{k^{\prime}}\int_{0}^{t}\dot{I}_{k^{\prime}}(t-\tau)M_{kk^{\prime}}\psi(\tau)\mathrm{d}\tau.$ (6)
Equation (6) extends Eq. (1) to complex networks. It is linear because of the assumption that the epidemic is in the initial stage and thus far from saturation. We assume exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$. Substituting this assumption into Eq. (6) we obtain
$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,.$ (7)
We call Eq. (7) the network Euler-Lotka equation (see section II in SI for an alternative derivation based on large deviation theory). Although the course of infectiousness and the network topology conspire in determining the spreading rate $\Lambda$, these two aspects appear as separated in Eq. (7). Indeed, the left hand side of Eq. (7) is uniquely determined by the network topology, while the right hand side only depends on $\psi(\tau)$ at given $\Lambda$. Equation (7) predicts that the epidemic threshold, i.e the condition $\Lambda=0$, is obtained by setting $R=1$, independently of $\psi(\tau)$.
We shall use Eq. (7) as an implicit relation for the unknown $\Lambda$. Using the first two moments of the generation-time distribution, we also obtain an explicit, although approximated expression:
$\Lambda\approx\frac{\ln R}{\langle\tau\rangle}+\left(\ln R\right)^{2}\frac{\sigma^{2}}{2\langle\tau\rangle^{3}}\,,$ (8)
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see section II-B in SI. An alternative is to combine Eq. (7) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to a lower bound on the exponential spreading rate of the epidemic:
$\Lambda\geq\frac{\ln R}{\langle\tau\rangle}\,.$ (9)
The bound in Eq. (9) is saturated when the epidemic spreads at regular time interval, i.e., when $\psi(\tau)$ is a Dirac delta function.
Synthetic networks
The network Euler-Lotka equation successfully predicts the rate of epidemic spreading on synthetic complex networks, Fig. 2-a and Fig. 2-b. Our battery of tests includes a broad range of models characterised by different degree distribution, assortativity, and clustering. For each model, we ran simulations using three different infectivity functions, see Fig. 2-c. The choice of the infectivity function qualitatively affect the behaviour of the epidemic. In particular, for peaked distributions, the exponential growth appears modulated by oscillations, since early generations of the epidemic are nearly synchronised. To extrapolate the leading exponential behaviour from these curves, we fit them to a generalised logistic function, which appear to capture well finite-size effects (see Methods). These fits also permits to define a finite-size parameter $\eta\in[0,1]$, which quantifies the relative impact of finite-size effects on the exponential spreading rate at short times (see Methods and section III in SI). The difference between the exponential spreading rate estimated from the simulations and from the network Euler-Lotka equation (7) is within $6\%$ error for all networks, see Fig. 2-f, except for the Barabási-Albert network with parameter $m=1$.
Figure 2: The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic networks.: Panel a: Average epidemic on a Erdős-Rényi network of size $N=10^{6}$ with $\overline{k}=4$ with Gamma distributed infection times with mean $7$ and variance $1$. The red curve is an average over 500 realisations of the epidemic. The exponential spreading rate (yellow dashed line) is obtained by fitting a generalised logistic function on the average trajectory (red dots). The black dashed line is the solution of the network Euler-Lotka equation with $R=R_{\mathrm{pert}}$ as given by Eq. (5). Panel b: Same as a, but with a Weibull distribution with shape $2$ and scale $2$ on a Watts-Strogatz network of size $N=10^{6}$. Panel c: Solution of the network Euler-Lotka equation compared with simulations for various networks. The procedure for a - b is repeated for various networks of sizes $10^{5}$ and $10^{6}$: Barabási-Albert (circles); Watts-Strogatz (diamonds); Erdős-Rényi (triangles); networks from the configuration model (CM: squares) with Lognormal and Borel degree distributions; networks from a generalisation of the configuration model (CCM: pentagons) whose assortativity and clustering coefficient have been modified (see details in section IV in SI). These simulations are repeated for three distributions: Weibull with parameter $2$ and scale $2$ (dashed green lines), Lognormal with mean $5$ and variance $3$ (solid blue lines), Gamma with mean $7$ and variance $1$ (dotted yellow lines). The black arrows point to the case of Barabási-Albert network for $m=1$. For these points, the values of the finite-size parameter are $\eta=0.68$ (Weibull), $\eta=0.245$ (Gamma), $\eta=0.400$ (Lognormal). In comparison, for $m=2$ we have $\eta=0.018$ (Weibull), $\eta=0.001$ (Gamma), $\eta=0.003$ (Lognormal). For networks other than Barabási-Albert networks, we find that on average $\eta=4.2\times 10^{-6}$ with worst case $\eta=1.2\times 10^{-5}$ (Lognormal network, Weibull distribution).
We now discuss the Barabási-Albert case in more detail. The parameter $m$ in the Barabási-Albert model represents the number of nodes a new nodes attaches to. In the case $m>1$, the network Euler-Lotka equation well predicts the spreading rate (Fig. 2f), despite these networks being scale-free, disassortative, and including loops. In contrast, the error in the estimate of the spreading rate for $m=1$ is approximately $15\%$. There are two possible causes for this fact. The first is that the exponential regime for $m=1$ is not clearly defined, as revealed by the finite-size parameter (see Fig. 2). The second is that, for $m=1$, the validity of the first-order perturbative expansion in $r$ might break down (see section V-A in SI). Indeed, a similar problem occurs when computing the average path length of the Barabási-Albert model for small $m$ [24]. This would also explain the large difference between $R_{\mathrm{pert}}$ and the leading eigenvalue of the reproductive matrix (see Table 1).
Figure 3: Epidemic spreading rate on real networks. We simulate epidemic spreading on real-world networks from ICON [25], SNAP [26] and KONECT [27]. All considered networks are undirected and without multiple links or self-connections. Their size varies from $N=10^{3}$ to $N=10^{6}$ (see section IV in SI). We simulate $500$ epidemics on each of these empirical networks with lognormally distributed infection times with mean $5$ and variance $3$. Panels a and b show the average trajectory for two different networks along with a representation of the actual network. Similarly to Figure 2, for each network, the exponential spreading rate (blue dashed-dotted line) is obtained by fitting a generalised logistic function to the average trajectory (red dots). The black dashed line represents the slope predicted by the network Euler-Lotka equation. Panel c: Exponential spreading rate for all empirical networks. The solid line is the solution of the Euler-Lotka equation, and the markers are the numerical fits. The colour of the marker denotes the value of the finite-size parameter $\eta$.
Empirical networks
We apply the Euler-Lotka equation to a large set of social, biological, technological, and transportation networks, see Fig. 3. These real-world networks exhibit more complex structures than network models, often characterised by tightly linked communities, strong degree-correlation and broad tails in the degree distribution. Therefore, they constitute a more challenging test. Surprisingly, the network Euler-Lotka equation holds well, see Fig. 3-a and Fig. 3-b for examples.
In general, we find that the prediction of the network Euler-Lotka equation well reproduces the fitted spreading rate whenever the finite-size parameter $\eta$ detects the presence of a clear exponential regime see Fig. 3-c. Despite some of these networks being highly clustered, using $R_{\mathrm{pert}}$ as expressed in Eq. (5) does not perform worse than $R$ (See tables section IV-E in SI). In Fig. 3-c, we have omitted three road networks (California, Texas, and Pennsylvania). For these networks, the generalized logistic fit fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories on these networks (see section IV-D in SI). Likely, the fact that these networks are embedded in a two-dimensional physical space prevents an exponential epidemic spread.
To better understand the cases where the error in the predicted spreading rate is large, we first scrutinise whether these networks are heavy-tailed. We quantify this by fitting the tail of the degree distribution by a power law of the form $k^{-\gamma}$ multiplied by a slowly varying function, see [28] and section VI in SI. We find large errors only when the finite-size parameter is large and $\gamma\leq 3$, see Fig. 4-a. In contrast, the error is weakly affected by the strength of clustering, see Fig. 4-b.
One problem with heavy-tailed networks is that the rapid infection of hubs depletes the tail of the degree distribution of the remaining nodes. This implies that, as the epidemic progresses, the degree distribution of the infected nodes does not remain invariant as assumed by our theory. To test this explanation, we study the average degree $\bar{k}_{n}$ of infected individuals in generation $n$. In a steady exponential regime, we expect this quantity to be independent of $n$, see Fig. 4-c. However, in scale-free networks, the average degree of a newly infected rapidly decays, see Fig. 4-d, leading to a non-steady spreading behaviour.
Figure 4: Finite-size effects impact early stages of an epidemic on scale-free networks . Panel a: Relative error between the fit and the Euler-Lotka equation as a function of the exponent $\gamma$ and the finite-size parameter $\eta$ (see Methods). The error is always low, except when the network is scale-free ($2<\gamma\leq 3$, between the dashed lines) and the finite-size parameter $\eta$ is large. Panel b: Error as a function of $\gamma$ and the exponent $\alpha$ of the degree-based clustering coefficient $c_{k}\sim(k-1)^{-\alpha}$ which can be interpreted as the strength of the clustering. Panel c: Average degree of individuals infected at the $n$th generation on a Erdős-Rényi graph of size $N=10^{6}$ and average degree $\overline{k}=4$. The dashed line is a power-law fit $\sim n^{-\delta}$ for $n\in[1,6]$. The fit yields $\delta\approx 0$, consistently with the presence of a clear exponential phase. Panel d: Average degree of individuals infected at the $n$th generation for the Youtube network, which has a diverging second moment ($\gamma\approx 2.2$ (see section VI in SI). The average degree decays as a power law with exponent $\delta\approx 4.48$, suggesting the absence of a clear exponential phase.
Discussion
In this paper, we developed a theory describing the exponential spreading rate of epidemics in populations arranged on a complex network. Our theory relates the spreading rate of an epidemic to the infectiousness and the social structure, while disentangling the impact of their distinct properties. Our approach accurately predicts the exponential rate of epidemic spreading on a wide range of model and real networks, aside from networks with heavy-tailed degree distributions where a clear exponential regime is absent.
When analysing real epidemics in their early stage, a departure from an exponential behaviour in the observed number of cases is often interpreted as a consequence of individual response or early containment measures (see, e.g., [29]). Our theory shows that non-exponential epidemic trajectories are relatively common in heavy-tailed networks, thus providing an alternative explanation for these observations.
We have focused for simplicity on undirected, static networks. The use of a static network is justified when the social structures evolve in a slower way than the time scale of the infection process. If however the time-scales are similar, the epidemic dynamics are more complex [30]. For example, bursty social interactions substantially affect epidemic spreading [31]. In principle, interactions could be also weighted, as duration and intensity of contacts also affect the probability of transmission [32]. The approach developed in this paper is flexible enough to be extended to these cases.
The theory developed in this paper has direct practical consequences for estimating the basic reproduction number $R$ [33]. Ideally, in the early stage of an epidemic, one could use the Euler-Lotka equation to estimate the basic reproduction number. However, this approach is fraught with practical difficulties. First, the curves in Fig. 2f show that a small error in the estimate of the exponential spreading rate $\Lambda$ can lead to a large uncertainty on $R$. Second, it is often difficult to reliably estimate the infectiousness distribution (see, e.g., [34, 35, 36]). The theory presented here permits to directly estimate $R$ from key properties of the interaction networks, such as moments of the degree distribution and clustering. These properties could be estimated, to some extent, based on contact tracing data. As a consequence, such approach could lead to improved estimates of the infectivity function as well.
Methods
Reproduction Number
We here derive the expression for the reproduction number, Eq. (5). We express the average number of infected individuals at generation $n\geq 1$ as
$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,,$ (10)
where the matrix $M_{kk^{\prime}}$ is given by Eq. (3). At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where $R$ is the leading eigenvalue of $M_{kk^{\prime}}$. For an arbitrary level of clustering and of degree correlation, the leading eigenvalue is not easy to compute exactly. We therefore approximate the eigenvalue by assuming weak clustering and treating the degree correlations as a perturbation of the uncorrelated case.
Degree correlations are summarised by the assortativity $r$ [37], defined as the Pearson correlation coefficient of the degrees of connected nodes:
$r=\frac{1}{\sigma_{r}^{2}}\sum_{kk^{\prime}}\left[kk^{\prime}P(k,k^{\prime})-\frac{k^{2}{k^{\prime}}^{2}}{\overline{k}^{2}}P(k)P(k^{\prime})\right],$ (11)
where $\sigma_{r}^{2}=\overline{k^{3}}/\overline{k}-\left(\overline{k^{2}}/\overline{k}\right)^{2}$ ensures that $-1\leq r\leq 1$ and $P(k,k^{\prime})=k^{\prime}P(k|k^{\prime})P(k^{\prime})/\overline{k}$ is the probability that a randomly chosen edge connects nodes of degrees $k$ and $k^{\prime}$.
We now analyse the clustering measure $m_{kk^{\prime}}$. While being a measure of clustering, $m_{kk^{\prime}}$ also contains information on the degree correlation of connected nodes. In contrast, the degree dependent clustering coefficient $c_{k}$ is independent of degree correlations [38]. This coefficient is related to $m_{kk^{\prime}}$ as $(k-1)c_{k}=\sum_{k^{\prime}}m_{kk^{\prime}}P(k^{\prime}|k)$. If we assume that $m_{kk^{\prime}}$ factorises into $k$ and $k^{\prime}$-dependent factors, we can express $m_{kk^{\prime}}$ as
$m_{kk^{\prime}}=\frac{(k-1)c_{k}(k^{\prime}-1)c_{k^{\prime}}}{m_{1}}$ (12)
where $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$. In particular, this factorisation is justified in the weak clustering limit [38]. Using this expression, we rewrite the reproduction matrix as
$M_{kk^{\prime}}=(k^{\prime}-1)\left[1-c(k^{\prime})\frac{(k-1)c_{k}}{m_{1}}\right]P(k|k^{\prime})\,$ (13)
To further simplify the problem, we neglect the dependence on $k$ of the probability that a node of degree $k^{\prime}$ connected to a node of degree $k$ is part of a triangle. This amounts to assume $(k-1)c_{k}/m_{1}\approx 1$, so that Eq. (13) becomes
$M_{kk^{\prime}}=(k^{\prime}-1)[1-c(k^{\prime})]P(k|k^{\prime})\,.$ (14)
We now treat $r$ as a small perturbation parameter and express the leading eigenvalue using perturbation theory. To this aim, we express the reproduction matrix as
$M_{kk^{\prime}}=M^{(r=0)}_{kk^{\prime}}+r\delta M_{kk^{\prime}},$ (15)
with
$M^{(r=0)}_{kk^{\prime}}$ $=(k^{\prime}-1)(1-c(k^{\prime}))\frac{k}{\overline{k}}P(k),$ (16)
$\delta M_{kk^{\prime}}$ $=(k^{\prime}-1)(1-c(k^{\prime}))\left(P^{(1)}(k|k^{\prime})-\frac{k}{\overline{k}}P(k)\right),$ (17)
where we have used that $P(k|k^{\prime})=kP(k)/\overline{k}$ for $r=0$ and $P^{(1)}(k|k^{\prime})$ is unspecified, but satisfies $\sum_{k}P^{(1)}(k|k^{\prime})=1$ and $\sum_{k}kP^{(1)}(k|k^{\prime})=k^{\prime}$. At this order in perturbation theory, the degrees of connected nodes are linearly correlated.
In the unperturbed case, the reproduction matrix has a single non-zero eigenvalue $R_{\mathrm{unper}}$ with left and right eigenvectors $u_{k}$ and $v_{k}$, normalised such that $\sum_{k}v_{k}u_{k}=1$:
$R_{\mathrm{unper}}$ $=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}-m_{1},$ (18)
$u_{k}$ $=\frac{(k-1)(1-c_{k})}{R_{\mathrm{unper}}}$ (19)
$v_{k}$ $=\frac{k}{\overline{k}}P(k).$ (20)
At the order of $\mathcal{O}(r)$, the leading eigenvalue is given by:
$R_{\mathrm{pert}}=R_{\mathrm{unper}}+r\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}+\mathcal{O}(r^{2}).$ (21)
$\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}$ $=\frac{1}{R_{\mathrm{unper}}}\sum_{kk^{\prime}}\left[(k-1)(1-c_{k})(k^{\prime}-1)(1-c(k^{\prime}))P^{(1)}(k|k^{\prime})P(k^{\prime})\frac{k^{\prime}}{\overline{k}}\right]-R_{\mathrm{unper}}.$ (22)
Expanding the terms involving clustering coefficients and evaluating them separately, we obtain
$R_{\mathrm{pert}}=R_{\mathrm{unper}}(1-r)+\frac{r}{R_{\mathrm{unper}}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)+\mathcal{O}(r^{2}),$ (23)
where we defined $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. At the first order, the right eigenvector associated to the unique leading eigenvalue is given by:
$v_{k}^{(1)}$ $=v_{k}+\frac{r}{R_{\mathrm{unper}}}\sum_{k^{\prime}}\delta M_{kk^{\prime}}v_{k^{\prime}}$
$\propto\left[1-r+\frac{r}{R_{\mathrm{unper}}}\left((k-1)-m_{1}\right)\right]kP(k)\,.$ (24)
As discussed, this eigenvector, if normalised, represents the relative probability that an infected individual has degree $k$ in the exponential regime.
Spreading rate of simulated epidemics: Generalised Logistic Function
In this section, we present our extrapolation of the exponential spreading rate $\Lambda$ of an epidemic from simulation data. In networks with heavy tails, finite-size effects alter the course of the epidemic at relatively early times. Likely for this reason, we have observed that functions such as a simple logistic regression fail to fit well the course of the epidemic. In contrast, we have found that the generalised logistic function
$I(t)=\frac{N}{\left(1+Qe^{-\Lambda\nu t}\right)^{1/\nu}},$ (25)
well fits the epidemic trajectories for the large majority of networks that we considered. In Eq. (25), $N$ the size of the network, $Q=(N/I_{0})^{\nu}-1$ and $I_{0}$ the initial number of infected individuals. The free parameters are $\Lambda$, $I_{0}$ and $\alpha$. Although each simulation starts with a single infected individual, we keep $I_{0}$ as a free parameter to better capture the behaviour of the epidemic at the early stages. We fit the logarithm of the generalised logistic function, and discard the first $\overline{k}$ infection events of each simulation, i.e. the times until $I(t)=\overline{k}$. The reason is that these infections should belong to the first generation $Z_{1}$, where the reproduction number is $\overline{k}$ rather than $R$.
A linearisation of $\log I(t)$ for small $t$ in Eq. (25) shows that:
$\log I(t)=\log I_{0}+\Lambda t\left(1-\left(\frac{I_{0}}{N}\right)^{\nu}\right)+\mathcal{O}(t^{2}).$ (26)
For $N\rightarrow\infty$, we obtain $I(t)\sim\exp(\Lambda t)$ as expected. Equation (26) leads us to define the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}$. For $\nu\to 1$, we recover a classic logistic function, while for $\nu\to 0$ the generalised logistic function tends to a Gompertz curve. For the synthetic and empirical networks that we studied, we have found that the finite-size parameter correlates with the curvature of $\log I(t)$ for small $t$, i.e. the deviation from an exponential behaviour (see section III in SI).
We call $\Lambda_{\rm FIT}$ the value of $\Lambda$ that makes Eq. (25) fit best the data. We call $\Lambda_{\rm EL}$ the value of $\Lambda$ obtained as solution of the Euler-Lotka equation. We measure the quality of our predictions by means of the symmetric relative error:
$\varepsilon=2\frac{\left|\Lambda_{\text{FIT}}-\Lambda_{\rm EL}\right|}{\Lambda_{\text{FIT}}+\Lambda_{\rm EL}}.$ (27)
The fits of all epidemic trajectories, together with tables presenting all values of the $\Lambda_{\rm FIT}$ and $\Lambda_{\rm EL}$ are available in section IV in SI.

### v2

@@
 The initial phase of an epidemic is often characterised by an exponential increase in the number of infected individuals.
-In well-mixed populations, this exponential increase is controlled by the basic reproduction number $R_{0}$ and the distribution of times between consecutive infection along an infection chain.
-However, we are still lacking a general understanding of how epidemics spread when individual interactions form a complex network.
+In a well-mixed population, this exponential increase is controlled by the basic reproduction number $R_{0}$ and the distribution of times between consecutive infection along an infection chain.
+However, real populations are characterised by a complex network of interactions, and are thus far from being well mixed.
 Here, we derive an expression for the rate of exponential spread of an epidemic on a complex network.
 We find that this rate is affected by the degree distribution, the network assortativity, and the level of clustering.
-Our result holds for a broad range of networks, aside from networks with very broad degree distribution, where no clear exponential regime is present.
+Our result holds for a broad range of networks, apart from networks with very broad degree distribution, where no clear exponential regime is present.
 The theory presented in this paper bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
-Epidemic modelling permits to rationalise evidences, test hypotheses, and inform disease control and other policy making.
-Epidemiological models should incorporate the key features that determine the epidemic spreading.
-One such feature is the fact that individuals are arranged in social structures [1, 2].
+Epidemic modelling permits rationalisation of evidences and can provide crucial information for policy making.
+Epidemiological models should incorporate the key features that determine epidemic spreading.
+One such feature is the arrangement of individuals in social structures [1].
 These social structures play a crucial role in the spreading of an epidemic, and in the effectiveness of intervention strategies [2, 3].
@@
 For example, at the early stages of COVID-19, the individuals with most contacts contributed disproportionately to the epidemic [4, 5].
-A convenient way of modelling social structures is by means of complex networks [7], in which nodes represent individuals and links signify that pairs of individuals are in close contact.
-A success of network theory has been to determine the epidemic threshold on networks [8], i.e., the transition point from a localised outbreak to a widespread epidemic.
-This theory has shown that networks whose degree distribution has a diverging variance due to heavy tails [9, 10] are extremely vulnerable to epidemics.
+A convenient way of modelling social structures is by means of complex networks [6], in which nodes represent individuals and links signify pairs of individuals that are in close contact.
+A success of network theory has been to determine the epidemic threshold on networks [7, 8, 9], i.e., the transition point from a localised outbreak to a widespread epidemic.
+This theory has shown that networks whose degree distribution has a diverging variance due to heavy tails [7, 10, 9] are extremely vulnerable to epidemics.
 A paradigmatic example is found in the Barabási-Albert model [11], where a network is constructed by preferentially attaching new nodes to existing nodes with high degree.
@@
 Many networks in nature possess both fat tails in their degree distribution and high level of clustering [12].
-A second key feature is the fact that the infectiousness of an individual, i.e., the rate at which an infected individuals infects its contacts, tends to change in time in a way that is particular to each disease.
-The underlying reason is that the infectiousness incorporates multiple evolving factors, such as the viral load of an individual, its social behaviour, and possibly environmental factors [13].
+Another key feature of an epidemic is its infectiousness, i.e., the rate at which an infected individual infects its contacts.
+The infectiousness incorporates multiple factors, such as the viral load of an individual, its social behaviour, and possibly environmental factors [13].
+Because of changes in these factors, the infectiousness of an individual tends to change in time in a way that is particular to each disease.
 This time-dependence profoundly impacts the way a disease spreads.
 Social structures and individual infectiousness are the key factors in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially.
-This stage is of central importance for novel epidemics, or during emergence of novel variants, often requiring rapid policy making.
-At later stages, other factors become relevant, such as individual recovery time, their behavioural change as a consequence of the epidemic [14], and the possibility of being reinfected.
-Despite the vast body of theory on epidemics in networks [10], results on the initial exponential regime are surprisingly sparse.
+This stage is of central importance for novel epidemics, or during emergence of novel variants, and it often requires rapid policy making.
+At later stages, other factors become relevant, such as individual recovery time, their behavioural change as a consequence of the epidemic [14], and the possible occurrence of reinfections.
+Despite the vast body of theory on epidemics in networks [9], general results on the speed of epidemic spreading are surprisingly sparse.
+The rate of epidemic spreading can be calculated for unclustered, uncorrelated networks, and with constant infectiousness [15].
 A recent theory estimates the average time it takes for an epidemic to reach an individual at a certain distance from patient zero [16].
@@
 Other approaches avoid a detailed description of social structures, and effectively model the heterogeneity of the population by assigning different risk propensities to individuals [18, 19].
-In this paper, we directly estimate the exponential rate of epidemic spreading on a complex network.
-Our approach is based on classic ideas of theoretical epidemiology, that we extend to complex networks.
+In this paper, we present a general theory that predicts the exponential rate of epidemic spreading on a complex network.
+In particular, we quantify the impact of key properties of the network, such as assortativity and clustering, on the reproduction number of the epidemic.
+We further link the reproduction number and the individual infectiousness with the epidemic spreading rate.
+We shall see that this approach leads to accurate predictions for a broad range of networks.
 Epidemic spreading in well-mixed populations
@@
 Substituting this assumption into Eq.
-(1) we obtain the Euler-Lotka equation
+(1) and taking the limit of large time, we obtain the Euler-Lotka equation
 $\frac{1}{R_{0}}=\langle e^{-\Lambda\tau}\rangle\,,$ (2)
@@
 a: Epidemic on a network.
-Nodes can be infected (red) or healthy (blue).
+Individuals (nodes) can be infected (red) or healthy (blue).
 Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbours, to whom they are connected via links.
@@
 Reproduction number on a complex network
-We define the basic reproduction number $R_{0}$ for a network as the average number of contacts of an infected individual in the early stage of the epidemic, excluding the one from which the infection came from.
+The basic reproduction number $R_{0}$ for a network is the average number of links of an infected individual in the early stage of the epidemic, excluding the one from which the infection came from.
 We also define the effective reproduction number $R$ as the average number of infections caused by an infected individual.
@@
 This permits us to identify $R$ as the leading eigenvalue of the reproduction matrix.
-The associated right eigenvector $v_{k^{\prime}}$ can be interpreted as the degree distribution of the infected individuals.
+The associated right eigenvector $v_{k^{\prime}}$ can be interpreted as the degree distribution of the infected individuals (see Methods).
 For large networks, diagonalising the reproduction matrix can be unwieldy.
-In the simple case where there is no clustering $(m_{kk^{\prime}}=0$ for all links) and the degrees of connected nodes are uncorrelated, we have
+In the simple case of an unclustered network $(m_{kk^{\prime}}=0$ for all links) in which the degrees of connected nodes are uncorrelated, we have
 $R=R_{0}=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}\,,$ (4)
@@
 We seek to extend it to the general case of clustered and correlated networks.
-We quantify degree correlations by the assortativity coefficient $r$.
+We quantify degree correlations by the assortativity coefficient $r$ (see Methods).
 When $r>0$, nodes with high degree tend to connect together whereas when $r<0$, they tend to connect to lower degree nodes.
-We find that, for small clustering coefficients and degree correlation, the basic reproduction number $R$ is approximated by
+We find that, for small clustering coefficients and degree correlation, the effective reproduction number $R$ is expressed by
 $R_{\mathrm{pert}}=(1-r)(R_{0}-m_{1})+\frac{r}{R_{0}-m_{1}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)\,,$ (5)
-where $c_{k}$ is the probability that two neighbours of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$ (see Methods).
-The reason for calling $R_{\mathrm{pert}}$ the expression in Eq.
-(5) is that this is obtained by a perturbative expansion in $r$ (see Methods).
+where $c_{k}$ is the probability that two neighbours of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$.
+Equation (5) is obtained by a perturbative expansion in $r$ (see Methods).
 We compare the expression for $R_{\mathrm{pert}}$ given in Eq.
@@
 We find that the relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the networks we considered, except for Barabási-Albert with $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$).
-We shall scrutinize the reason for this discrepancy later in the paper.
+We shall analyse the reason for this discrepancy later in the paper.
+Approximating $R$ with $R_{0}$ given in Eq.
+(4) leads to substantially larger errors, supporting that clustering and assortativity need to be taken into account to appropriately estimate $R$.
 Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq.
+(5) and $R_{0}$ using Eq.
 (4) for a given network, we always interpret the averages as empirical averages.
-Because of this choice, moments of the degree distribution are always finite, even in very heterogeneous networks.
+With this choice, moments of the degree distribution are always finite, even in very heterogeneous networks.
 In practical cases, Eq.
@@
 (5) is usually rather small, except when the degree distribution has heavy tails.
-In contrast, clustering systematically lowers $R_{\mathrm{pert}}$, regardless of the degree distribution.
-Therefore, approximating the reproduction number with $R_{0}$ leads to an overestimate in the presence of clustering.
+In contrast, clustering systematically reduces $R_{\mathrm{pert}}$, regardless of the degree distribution.
+For this reason, we find that $R_{0}$ overestimates $R$ in all the synthetic networks we considered.
 Table 1: Effect of clustering and assortativity on the reproduction number.
-Network Assortativity $r$ Clustering $\overline{c}$ $R_{0}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ $R_{\mathrm{pert}}$ $R$
-$\text{BA}_{1}$0.011 0 13.9 13.9 6.6 6.8 8.5
-$\text{BA}_{2}$0.008 0.0001 20.4 20.3 14.2 14.4 15.5
-$\text{BA}_{12}$0.004 0.0003 85.4 83.6 80.4 79.5 80.8
-ER6 0.001 0.000 6.0 6.0 6.0 6.0 6.0
-[…]
-WS8,2 -0.022 0.334 7.2 4.9 7.0 4.9 4.9
-WS6,2 -0.029 0.315 5.2 3.6 5.0 3.6 3.6
-$\text{LN}_{4,1}\;\alpha=0.5$ 0.007 0.161 3.7 3.2 3.8 3.2 3.2
-Parameters associated with the epidemic spreading for different models of complex networks: BAm: Barabási-Albert model with parameter $m$.
+Network $R_{0}$ $R$ $R_{\mathrm{pert}}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ Assortativity $r$ Clustering $\overline{c}$
+$\text{BA}_{1}$ 13.9 8.5 6.8 13.9 6.6 -0.011 0
+$\text{BA}_{2}$ 20.4 15.5 14.4 20.3 14.2 -0.008 0.0001
+$\text{BA}_{12}$ 85.4 80.8 79.5 83.6 80.4 -0.004 0.0003
+$\text{ER}_{6}$ 6.0 6.0 6.0 6.0 6.0 0.001 0.000
+$\text{ER}_{10}\;\alpha=0.2$ 10.0 9.0 9.1 9.0 10.7 0.062 0.108
+$\text{WS}_{8,2}$ 7.2 4.9 4.9 4.9 7.0 -0.022 0.334
+$\text{WS}_{6,2}$ 5.2 3.6 3.6 3.6 5.0 -0.029 0.315
+$\text{LN}_{4,1}\;\alpha=0.5$ 3.7 3.2 3.2 3.2 3.8 0.007 0.161
+Parameters associated with the epidemic spreading for different models of complex networks.
+The basic reproduction number $R_{0}$ is given by Eq.
+(4).
+The effective reproduction number $R$ is the numerically computed leading eigenvalue of Eq.
+(3).
+We compare it with the analytical expression of $R_{\mathrm{pert}}$ given in Eq.
+(5).
+To assess the impact of degree correlations and clustering, we also compute $R_{\mathrm{pert}}$ for $r=0$ and for $\overline{c}=0$, where we have defined $\bar{c}=\sum_{k}c_{k}P(k)$.
+BAm: Barabási-Albert model with parameter $m$.
 $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$.
@@
 $\text{ER}_{10}\;\alpha=0.2$: A Erdős-Rényi network in which we imposed clustering using the same algorithm.
-We have defined the average clustering coefficient $\bar{c}=\sum_{k}c_{k}P(k)$.
-The reproduction number $R$ is the numerically computed leading eigenvalue of Eq.
-(3), whereas we obtain $R_{\mathrm{pert}}$ by evaluating Eq.
-(5).
-To assess the impact of degree correlations and clustering, we also compute $R_{\mathrm{pert}}$ for $\overline{c}=0$.
 Network Euler-Lotka equation
@@
 Substituting this assumption into Eq.
-(6) we obtain
-$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,.$ (7)
-We call Eq.
-(7) the network Euler-Lotka equation (see section II in SI for an alternative derivation based on large deviation theory).
-Although the course of infectiousness and the network topology conspire in determining the spreading rate $\Lambda$, these two aspects appear as separated in Eq.
-(7).
-Indeed, the left hand side of Eq.
-(7) is uniquely determined by the network topology, while the right hand side only depends on $\psi(\tau)$ at given $\Lambda$.
-Equation (7) predicts that the epidemic threshold, i.e the condition $\Lambda=0$, is obtained by setting $R=1$, independently of $\psi(\tau)$.
-We shall use Eq. (7) as an implicit relation for the unknown $\Lambda$. Using the first two moments of the generation-time distribution, we also obtain an explicit, although approximated expression:
+(6) we obtain, for large time,
+$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,,$ (7)
+[…]
+In general, we shall use Eq.
+(7) as an implicit relation for the unknown $\Lambda$.
+It is possible to explicitly express $\Lambda$ from Eq.
+(7), but only in an approximate way.
+In particular, using the first two moments of the generation-time distribution, we obtain
 $\Lambda\approx\frac{\ln R}{\langle\tau\rangle}+\left(\ln R\right)^{2}\frac{\sigma^{2}}{2\langle\tau\rangle^{3}}\,,$ (8)
-where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see section II-B in SI.
+where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see Section LABEL:SI-gaussian in SI.
 An alternative is to combine Eq.
@@
 The choice of the infectivity function qualitatively affect the behaviour of the epidemic.
-In particular, for peaked distributions, the exponential growth appears modulated by oscillations, since early generations of the epidemic are nearly synchronised.
-To extrapolate the leading exponential behaviour from these curves, we fit them to a generalised logistic function, which appear to capture well finite-size effects (see Methods).
-These fits also permits to define a finite-size parameter $\eta\in[0,1]$, which quantifies the relative impact of finite-size effects on the exponential spreading rate at short times (see Methods and section III in SI).
-The difference between the exponential spreading rate estimated from the simulations and from the network Euler-Lotka equation (7) is within $6\%$ error for all networks, see Fig.
-2-f, except for the Barabási-Albert network with parameter $m=1$.
-Figure 2: The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic networks.: Panel a: Average epidemic on a Erdős-Rényi network of size $N=10^{6}$ with $\overline{k}=4$ with Gamma distributed infection times with mean $7$ and variance $1$.
+In particular, for peaked distributions, the exponential growth appears modulated by oscillations (Fig.
+2-a), since early generations of the epidemic are nearly synchronised.
+To extrapolate the leading exponential behaviour from these curves, we fit them to a generalised logistic function, which appear to well capture finite-size effects (see Methods).
+The fits also permits to define a finite-size parameter $\eta\in[0,1]$, which quantifies the relative impact of finite-size effects on the exponential spreading rate at short times (see Methods and Section LABEL:SI-sec:eta in SI).
+When predicting the exponential growth rate using the network Euler-Lotka equation, we use $R_{\text{pert}}$ as defined in Eq.
+(5) as the reproduction number.
+As expected, the predictions are considerably worse if we instead use $R_{0}$, see Section LABEL:SI-sec:R_vs_R0 in SI.
+The difference between the exponential spreading rate estimated from the simulations and from the network Euler-Lotka equation (7) is within $6\%$ for all networks, see Fig.
+2-c.
+The only exception is the Barabási-Albert network with parameter $m=1$ (average error $15\%$), where $m$ represents the number of nodes a new node attaches to.
+Figure 2: The network Euler-Lotka equation predicts the rate of epidemic spreading on synthetic networks.: Panel a: Average epidemic on a Erdős-Rényi network of size $N=10^{6}$ with $\overline{k}=4$ with Gamma distributed infection times.
 The red curve is an average over 500 realisations of the epidemic.
-The exponential spreading rate (yellow dashed line) is obtained by fitting a generalised logistic function on the average trajectory (red dots).
-The black dashed line is the solution of the network Euler-Lotka equation with $R=R_{\mathrm{pert}}$ as given by Eq.
+We estimate the exponential spreading rate (yellow dashed line) by fitting a generalised logistic function on the average trajectory (red dots), see Methods.
+The black dashed line is the solution of the network Euler-Lotka Equation (7) with $R=R_{\mathrm{pert}}$ as given by Eq.
 (5).
-Panel b: Same as a, but with a Weibull distribution with shape $2$ and scale $2$ on a Watts-Strogatz network of size $N=10^{6}$.
+Panel b: Same as a, but with a Weibull distribution on a Watts-Strogatz network of size $N=10^{6}$.
 Panel c: Solution of the network Euler-Lotka equation compared with simulations for various networks.
-The procedure for a - b is repeated for various networks of sizes $10^{5}$ and $10^{6}$: Barabási-Albert (circles); Watts-Strogatz (diamonds); Erdős-Rényi (triangles); networks from the configuration model (CM: squares) with Lognormal and Borel degree distributions; networks from a generalisation of the configuration model (CCM: pentagons) whose assortativity and clustering coefficient have been modified (see details in section IV in SI).
-These simulations are repeated for three distributions: Weibull with parameter $2$ and scale $2$ (dashed green lines), Lognormal with mean $5$ and variance $3$ (solid blue lines), Gamma with mean $7$ and variance $1$ (dotted yellow lines).
-The black arrows point to the case of Barabási-Albert network for $m=1$.
-For these points, the values of the finite-size parameter are $\eta=0.68$ (Weibull), $\eta=0.245$ (Gamma), $\eta=0.400$ (Lognormal).
-In comparison, for $m=2$ we have $\eta=0.018$ (Weibull), $\eta=0.001$ (Gamma), $\eta=0.003$ (Lognormal).
-For networks other than Barabási-Albert networks, we find that on average $\eta=4.2\times 10^{-6}$ with worst case $\eta=1.2\times 10^{-5}$ (Lognormal network, Weibull distribution).
-We now discuss the Barabási-Albert case in more detail.
-The parameter $m$ in the Barabási-Albert model represents the number of nodes a new nodes attaches to.
-In the case $m>1$, the network Euler-Lotka equation well predicts the spreading rate (Fig.
-2f), despite these networks being scale-free, disassortative, and including loops.
-In contrast, the error in the estimate of the spreading rate for $m=1$ is approximately $15\%$.
-There are two possible causes for this fact.
-The first is that the exponential regime for $m=1$ is not clearly defined, as revealed by the finite-size parameter (see Fig.
+The procedure for a - b is repeated for various networks of sizes $10^{5}$ and $10^{6}$: Barabási-Albert (circles); Watts-Strogatz (diamonds); Erdős-Rényi (triangles); networks from the configuration model (CM: squares) with Lognormal and Borel degree distributions; networks from a generalisation of the configuration model (CCM: pentagons) whose assortativity and clustering coefficient have been modified (see Section LABEL:SI-sec:simulations in SI).
+These simulations are repeated for three generation time distributions: Weibull (dashed green lines), Lognormal (solid blue lines), Gamma (dotted yellow lines).
+The expressions of these distributions are given in SI Section LABEL:SI-sec:simulations.
+There are two causes for this exception.
+The first is that the exponential regime for this model is not clearly defined, as revealed by the finite-size parameter (see Fig.
 2).
-The second is that, for $m=1$, the validity of the first-order perturbative expansion in $r$ might break down (see section V-A in SI).
-Indeed, a similar problem occurs when computing the average path length of the Barabási-Albert model for small $m$ [24].
-This would also explain the large difference between $R_{\mathrm{pert}}$ and the leading eigenvalue of the reproductive matrix (see Table 1).
-Figure 3: Epidemic spreading rate on real networks.
+This issue is likely related with the fact that the second moment $\bar{k^{2}}$ of the degree distribution diverges in the infinite size limit.
+The second cause is that, for $m=1$, the first-order perturbative expansion in $r$ loses accuracy, as revealed by the difference between $R_{\mathrm{pert}}$ and the leading eigenvalue of the reproductive matrix (see Table 1).
+Figure 3: The network Euler-Lotka equation predicts the epidemic spreading rate on real-world networks.
 We simulate epidemic spreading on real-world networks from ICON [25], SNAP [26] and KONECT [27].
 All considered networks are undirected and without multiple links or self-connections.
-Their size varies from $N=10^{3}$ to $N=10^{6}$ (see section IV in SI).
-We simulate $500$ epidemics on each of these empirical networks with lognormally distributed infection times with mean $5$ and variance $3$.
+Their size varies from $N=10^{3}$ to $N=10^{6}$ (see Section LABEL:SI-sec:real_networks in SI).
+We simulate $500$ epidemics on each of these real-world networks with lognormally distributed infection times, see SI Section LABEL:SI-sec:simulations.
 Panels a and b show the average trajectory for two different networks along with a representation of the actual network.
-Similarly to Figure 2, for each network, the exponential spreading rate (blue dashed-dotted line) is obtained by fitting a generalised logistic function to the average trajectory (red dots).
+Similarly to Figure 2, for each network, we obtain the exponential spreading rate (blue dashed-dotted line) by fitting a generalised logistic function to the average trajectory (red dots).
 The black dashed line represents the slope predicted by the network Euler-Lotka equation.
-Panel c: Exponential spreading rate for all empirical networks.
-The solid line is the solution of the Euler-Lotka equation, and the markers are the numerical fits.
+Panel c: Exponential spreading rate for all real-world networks.
+The solid line is the solution of the Euler-Lotka equation.
+The markers are the numerical fits.
 The colour of the marker denotes the value of the finite-size parameter $\eta$.
-Empirical networks
-We apply the Euler-Lotka equation to a large set of social, biological, technological, and transportation networks, see Fig.
+Real-world networks
+We apply the network Euler-Lotka equation to a large set of social, biological, technological, and transportation networks, see Fig.
 3.
-These real-world networks exhibit more complex structures than network models, often characterised by tightly linked communities, strong degree-correlation and broad tails in the degree distribution.
+These real-world networks exhibit more complex structures than synthetic networks.
+They are often characterised by tightly linked communities, strong degree-correlation and broad tails in the degree distribution.
 Therefore, they constitute a more challenging test.
-Surprisingly, the network Euler-Lotka equation holds well, see Fig.
+Nevertheless, the network Euler-Lotka equation holds well, see Fig.
 3-a and Fig.
 3-b for examples.
-In general, we find that the prediction of the network Euler-Lotka equation well reproduces the fitted spreading rate whenever the finite-size parameter $\eta$ detects the presence of a clear exponential regime see Fig.
+More generally, we find that the prediction of the network Euler-Lotka equation well reproduces the fitted spreading rate whenever the finite-size parameter $\eta$ detects a clear exponential regime, see Fig.
 3-c.
-Despite some of these networks being highly clustered, using $R_{\mathrm{pert}}$ as expressed in Eq.
-(5) does not perform worse than $R$ (See tables section IV-E in SI).
+Although some of these networks are highly clustered, using $R_{\mathrm{pert}}$ as expressed by Eq.
+(5) does not perform worse than $R$ (See Section LABEL:SI-sec:R_vs_R0 in SI).
 In Fig.
 3-c, we have omitted three road networks (California, Texas, and Pennsylvania).
-For these networks, the generalized logistic fit fails to converge.
-However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories on these networks (see section IV-D in SI).
+For these networks, the generalised logistic fit fails to converge.
+However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories on these networks (see Section LABEL:SI-power-law-fit in SI).
 Likely, the fact that these networks are embedded in a two-dimensional physical space prevents an exponential epidemic spread.
-To better understand the cases where the error in the predicted spreading rate is large, we first scrutinise whether these networks are heavy-tailed.
-We quantify this by fitting the tail of the degree distribution by a power law of the form $k^{-\gamma}$ multiplied by a slowly varying function, see [28] and section VI in SI.
-We find large errors only when the finite-size parameter is large and $\gamma\leq 3$, see Fig.
+For networks in which our fit detects small finite-size effects ($\eta<0.5$), the predictions of the network Euler-Lotka equation using $R_{\text{pert}}$ are more accurate than those using $R_{0}$.
+However, for larger values of $\eta$, $R_{0}$ performs slightly better on average (see Section LABEL:SI-sec:R_vs_R0 in SI).
+Even in these cases, $R_{\mathrm{pert}}$ provides a better approximation of the reproduction number $R$, see SI Fig.
+LABEL:SI-fig:Rpert_vs_eta-b.
+This suggests that the advantage of $R_{0}$ for large $\eta$ might be due to a compensatory effect: using $R_{0}$ tends to overestimate the reproduction number, while the network Euler-Lotka equation tends to underestimate $\Lambda$ compared to $\Lambda_{\mathrm{FIT}}$.
+We aim at understanding why, for certain real-world networks, the predicted spreading rate deviates from the fitted one.
+We find large errors only for scale-free networks in which the finite-size parameter $\eta$ is also large, see Fig.
 4-a.
+We call a network scale-free if the tail of its degree distribution are fitted by a power law of the form $k^{-\gamma}$, with $\gamma\leq 3$, multiplied by a slowly varying function, see [28] and Section LABEL:SI-sec:tail in SI.
 In contrast, the error is weakly affected by the strength of clustering, see Fig.
 4-b.
-One problem with heavy-tailed networks is that the rapid infection of hubs depletes the tail of the degree distribution of the remaining nodes.
+One issue with heavy-tailed networks is that the rapid infection of hubs depletes the tail of the degree distribution of the remaining nodes.
 This implies that, as the epidemic progresses, the degree distribution of the infected nodes does not remain invariant as assumed by our theory.
@@
 4-c.
-However, in scale-free networks, the average degree of a newly infected rapidly decays, see Fig.
-4-d, leading to a non-steady spreading behaviour.
+However, in scale-free networks, the average degree of infected individuals appears to decay as a power-law of their generation, see Fig.
+4-d, leading to a non-steady spreading.
 Figure 4: Finite-size effects impact early stages of an epidemic on scale-free networks .
 Panel a: Relative error between the fit and the Euler-Lotka equation as a function of the exponent $\gamma$ and the finite-size parameter $\eta$ (see Methods).
-The error is always low, except when the network is scale-free ($2<\gamma\leq 3$, between the dashed lines) and the finite-size parameter $\eta$ is large.
+The error is always low, except when the network is scale-free ($2<\gamma\leq 3$, region between the two dashed lines) and the finite-size parameter $\eta$ is large.
 Panel b: Error as a function of $\gamma$ and the exponent $\alpha$ of the degree-based clustering coefficient $c_{k}\sim(k-1)^{-\alpha}$ which can be interpreted as the strength of the clustering.
@@
 The dashed line is a power-law fit $\sim n^{-\delta}$ for $n\in[1,6]$.
-The fit yields $\delta\approx 0$, consistently with the presence of a clear exponential phase.
-Panel d: Average degree of individuals infected at the $n$th generation for the Youtube network, which has a diverging second moment ($\gamma\approx 2.2$ (see section VI in SI).
-The average degree decays as a power law with exponent $\delta\approx 4.48$, suggesting the absence of a clear exponential phase.
+The fit yields $\delta\approx 0$, consistently with a steady exponential spreading.
+Panel d: Average degree of individuals infected at the $n$th generation for the Youtube network, which is scale-free ($\gamma\approx 2.2$ (see Section LABEL:SI-sec:tail in SI).
+The average degree decays as a power law with exponent $\delta\approx 4.48$ and is therefore not steady.
 Discussion
 In this paper, we developed a theory describing the exponential spreading rate of epidemics in populations arranged on a complex network.
-Our theory relates the spreading rate of an epidemic to the infectiousness and the social structure, while disentangling the impact of their distinct properties.
-Our approach accurately predicts the exponential rate of epidemic spreading on a wide range of model and real networks, aside from networks with heavy-tailed degree distributions where a clear exponential regime is absent.
-When analysing real epidemics in their early stage, a departure from an exponential behaviour in the observed number of cases is often interpreted as a consequence of individual response or early containment measures (see, e.g., [29]).
-Our theory shows that non-exponential epidemic trajectories are relatively common in heavy-tailed networks, thus providing an alternative explanation for these observations.
-We have focused for simplicity on undirected, static networks.
-The use of a static network is justified when the social structures evolve in a slower way than the time scale of the infection process.
-If however the time-scales are similar, the epidemic dynamics are more complex [30].
-For example, bursty social interactions substantially affect epidemic spreading [31].
-In principle, interactions could be also weighted, as duration and intensity of contacts also affect the probability of transmission [32].
-The approach developed in this paper is flexible enough to be extended to these cases.
-The theory developed in this paper has direct practical consequences for estimating the basic reproduction number $R$ [33].
-Ideally, in the early stage of an epidemic, one could use the Euler-Lotka equation to estimate the basic reproduction number.
+Our theory relates the spreading rate to the infectiousness and the social structure, while disentangling the impact of their distinct properties.
+Our approach accurately predicts the spreading rate on a wide range of model and real networks, aside from networks with heavy-tailed degree distributions where the exponential regime is hindered by finite-size effects.
+When analysing the early stage of real epidemics, a departure from an exponential behaviour in the observed number of cases is often interpreted as a consequence of individual response or early containment measures (see, e.g., [29]).
+Our results show that non-exponential epidemic trajectories are common in heavy-tailed networks, thus providing an alternative explanation for these observations.
+The theory developed in this paper can be also used to estimate the basic reproduction number $R$ [30].
+Ideally, in the early stage of an epidemic, one could use the Euler-Lotka equation to deduce $R$.
 However, this approach is fraught with practical difficulties.
@@
 2f show that a small error in the estimate of the exponential spreading rate $\Lambda$ can lead to a large uncertainty on $R$.
-Second, it is often difficult to reliably estimate the infectiousness distribution (see, e.g., [34, 35, 36]).
-The theory presented here permits to directly estimate $R$ from key properties of the interaction networks, such as moments of the degree distribution and clustering.
-These properties could be estimated, to some extent, based on contact tracing data.
+Second, it is often difficult to reliably estimate the infectiousness distribution (see, e.g., [31, 32, 33] for the case of COVID-19).
+Instead, our theory permits to directly estimate $R$ from structural properties of the interaction networks, such as moments of the degree distribution and clustering.
+These properties could be inferred, to some extent, based on contact tracing data [34].
 As a consequence, such approach could lead to improved estimates of the infectivity function as well.
+We have focused for simplicity on static networks, which is justified when the social structures evolve in a slower way than the time scale of the infection process.
+If however the time-scales are similar, epidemic dynamics are more complex [35].
+For example, bursty social interactions substantially affect epidemic spreading [36].
+In principle, interactions could be also weighted, as duration and intensity of contacts affect the probability of transmission [37].
+The approach developed in this paper is flexible enough to be extended to these scenarios.
 Methods
 Reproduction Number
-We here derive the expression for the reproduction number, Eq. (5). We express the average number of infected individuals at generation $n\geq 1$ as
-$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,,$ (10)
-where the matrix $M_{kk^{\prime}}$ is given by Eq.
+We here derive the expression for the reproduction number, Eq.
 (5).
+We express the average number of infected individuals at generation $n\geq 1$ in terms of the reproduction matrix introduced in Eq.
+(3) by summing over the degree of individuals at each generation:
+$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,.$ (10)
 At the first generation, we obtain $Z_{1}=\bar{k}$.
@@
 To further simplify the problem, we neglect the dependence on $k$ of the probability that a node of degree $k^{\prime}$ connected to a node of degree $k$ is part of a triangle.
-This amounts to assume $(k-1)c_{k}/m_{1}\approx 1$, so that Eq.
+This amounts to assume $(k-1)c_{k}/m_{1}=1$, so that Eq.
 (13) becomes
@@
 At this order in perturbation theory, the degrees of connected nodes are linearly correlated.
-In the unperturbed case, the reproduction matrix has a single non-zero eigenvalue $R_{\mathrm{unper}}$ with left and right eigenvectors $u_{k}$ and $v_{k}$, normalised such that $\sum_{k}v_{k}u_{k}=1$:
-$R_{\mathrm{unper}}$ $=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}-m_{1},$ (18)
-$u_{k}$ $=\frac{(k-1)(1-c_{k})}{R_{\mathrm{unper}}}$ (19)
-$v_{k}$ $=\frac{k}{\overline{k}}P(k).$ (20)
-At the order of $\mathcal{O}(r)$, the leading eigenvalue is given by:
-$R_{\mathrm{pert}}=R_{\mathrm{unper}}+r\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}+\mathcal{O}(r^{2}).$ (21)
-$\sum_{kk^{\prime}}u_{k}\delta M_{kk^{\prime}}v_{k^{\prime}}$ $=\frac{1}{R_{\mathrm{unper}}}\sum_{kk^{\prime}}\left[(k-1)(1-c_{k})(k^{\prime}-1)(1-c(k^{\prime}))P^{(1)}(k|k^{\prime})P(k^{\prime})\frac{k^{\prime}}{\overline{k}}\right]-R_{\mathrm{unper}}.$ (22)
-Expanding the terms involving clustering coefficients and evaluating them separately, we obtain
-$R_{\mathrm{pert}}=R_{\mathrm{unper}}(1-r)+\frac{r}{R_{\mathrm{unper}}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)+\mathcal{O}(r^{2}),$ (23)
-where we defined $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. At the first order, the right eigenvector associated to the unique leading eigenvalue is given by:
-$v_{k}^{(1)}$ $=v_{k}+\frac{r}{R_{\mathrm{unper}}}\sum_{k^{\prime}}\delta M_{kk^{\prime}}v_{k^{\prime}}$
-$\propto\left[1-r+\frac{r}{R_{\mathrm{unper}}}\left((k-1)-m_{1}\right)\right]kP(k)\,.$ (24)
-As discussed, this eigenvector, if normalised, represents the relative probability that an infected individual has degree $k$ in the exponential regime.
+At the first order in $r$, we obtain
+$R_{\mathrm{pert}}=(R_{0}-m_{1})(1-r)+\frac{r}{R_{0}-m_{1}}\left(\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}\right)+\mathcal{O}(r^{2})\,,$ (18)
+see SI, where we defined $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. At this order, the right eigenvector associated to the unique leading eigenvalue is given by:
+$v_{k}$ $=\frac{kP(k)}{\bar{k}}+\frac{r}{R_{0}-m_{1}}\sum_{k^{\prime}}\delta M_{kk^{\prime}}\frac{k^{\prime}P(k^{\prime})}{\bar{k}}$
+$\propto\left[1-r+\frac{r}{R_{0}-m_{1}}\left((k-1)-m_{1}\right)\right]kP(k)\,.$ (19)
+This eigenvector, if normalised, represents the relative probability that an infected individual has degree $k$ in the exponential regime.
 Spreading rate of simulated epidemics: Generalised Logistic Function
-In this section, we present our extrapolation of the exponential spreading rate $\Lambda$ of an epidemic from simulation data.
+We describe our extrapolation of the exponential spreading rate $\Lambda$ of an epidemic from simulation data.
 In networks with heavy tails, finite-size effects alter the course of the epidemic at relatively early times.
@@
 In contrast, we have found that the generalised logistic function
-$I(t)=\frac{N}{\left(1+Qe^{-\Lambda\nu t}\right)^{1/\nu}},$ (25)
+$I(t)=\frac{N}{\left(1+Qe^{-\Lambda_{0}\nu t}\right)^{1/\nu}},$ (20)
 well fits the epidemic trajectories for the large majority of networks that we considered.
 In Eq.
-(25), $N$ the size of the network, $Q=(N/I_{0})^{\nu}-1$ and $I_{0}$ the initial number of infected individuals.
-The free parameters are $\Lambda$, $I_{0}$ and $\alpha$.
-Although each simulation starts with a single infected individual, we keep $I_{0}$ as a free parameter to better capture the behaviour of the epidemic at the early stages.
-We fit the logarithm of the generalised logistic function, and discard the first $\overline{k}$ infection events of each simulation, i.e.
-the times until $I(t)=\overline{k}$.
-The reason is that these infections should belong to the first generation $Z_{1}$, where the reproduction number is $\overline{k}$ rather than $R$.
+(20), $N$ the size of the network and $Q=(N/I_{0})^{\nu}-1$ where $I_{0}$ the initial number of infected individuals.
+Although each simulation starts with a single infected individual, we consider $I_{0}$ as a free parameter to better capture the behaviour of the epidemic at the early stages.
+The remaining free parameters are $\Lambda_{0}$ and $\alpha$.
+We fit the logarithm of the generalised logistic function, and discard the first $\overline{k}$ infection events of each simulation, i.e., the times until $I(t)=\overline{k}$.
+The reason is that these infections should belong to the first generation $Z_{1}$, in which the average number of secondary infections per individual is $\overline{k}$ rather than $R$.
 A linearisation of $\log I(t)$ for small $t$ in Eq. (20) shows that:
-$\log I(t)=\log I_{0}+\Lambda t\left(1-\left(\frac{I_{0}}{N}\right)^{\nu}\right)+\mathcal{O}(t^{2}).$ (26)
-For $N\rightarrow\infty$, we obtain $I(t)\sim\exp(\Lambda t)$ as expected.
-Equation (26) leads us to define the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}$.
+$\log I(t)=\log I_{0}+\Lambda_{0}t\left(1-\left(\frac{I_{0}}{N}\right)^{\nu}\right)+\mathcal{O}(t^{2}).$ (21)
+For $N\rightarrow\infty$, we obtain $I(t)\sim\exp(\Lambda_{0}t)$.
+Equation (21) leads us to define the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}$ and the finite-size growth rate $\Lambda=\Lambda_{0}(1-\eta)$.
 For $\nu\to 1$, we recover a classic logistic function, while for $\nu\to 0$ the generalised logistic function tends to a Gompertz curve.
-For the synthetic and empirical networks that we studied, we have found that the finite-size parameter correlates with the curvature of $\log I(t)$ for small $t$, i.e.
-the deviation from an exponential behaviour (see section III in SI).
+For the synthetic and real-world networks that we studied, we have found that the finite-size parameter correlates with the curvature of $\log I(t)$ for small $t$, i.e.
+the deviation from an exponential behaviour (see Section LABEL:SI-sec:eta in SI).
 We call $\Lambda_{\rm FIT}$ the value of $\Lambda$ that makes Eq.
@@
 $\varepsilon=2\frac{\left|\Lambda_{\text{FIT}}-\Lambda_{\rm EL}\right|}{\Lambda_{\text{FIT}}+\Lambda_{\rm EL}}.$ (22)
-The fits of all epidemic trajectories, together with tables presenting all values of the $\Lambda_{\rm FIT}$ and $\Lambda_{\rm EL}$ are available in section IV in SI.
+The fits of all epidemic trajectories, together with tables presenting all values of the $\Lambda_{\rm FIT}$ and $\Lambda_{\rm EL}$ are presented in Section LABEL:SI-sec:simulations in SI.
+Acknowledgements
+We thank C.
+Castellano and I.
+Neri for comments on a preliminary version of this manuscript.
+We thank Joshua Weitz for discussions.
+We are grateful for the help and support provided by the Scientific Computing and Data Analysis section of Core Facilities at OIST.